# Unveiling quantum phase transitions from traps in variational quantum algorithms

### v1

@@ Abstract
Understanding quantum phase transitions in physical systems is fundamental to characterize their behaviour at small temperatures. Achieving this requires both accessing good approximations to the ground state and identifying order parameters to distinguish different phases. Addressing these challenges, our work introduces a hybrid algorithm that combines quantum optimization with classical machine learning. This approach leverages the capability of near-term quantum computers to prepare locally trapped states through finite optimization. Specifically, we utilize LASSO for identifying conventional phase transitions and the Transformer model for topological transitions, applying these with a sliding window of Hamiltonian parameters to learn appropriate order parameters and estimate the critical points accurately. We verified the effectiveness of our method with numerical simulation and real-hardware experiments on Rigetti’s Ankaa 9Q-1 quantum computer. Our protocol not only provides a robust framework for investigating quantum phase transitions using shallow quantum circuits but also significantly enhances efficiency and precision, opening new avenues in the integration of quantum computing and machine learning.
@@ I Introduction
Quantum phase diagrams are a fundamental tool to characterize the behavior of quantum systems under various external conditions, such as temperature and magnetic fields [1, 2, 3, 4, 5, 6]. A better understanding of quantum phase transitions finds application in the field of material science, where it can aid in the development of novel materials with unique properties. For instance, understanding superconductor–insulator phase transitions is instrumental in the research and development of high-temperature superconductors [7, 8]. Due to the intrinsic complexity of quantum systems, the accurate determination of the ground state phase diagram is a formidable challenge [9, 10]. Experimentally achieving and maintaining the ground state is also difficult, requiring precise control of conditions and often near-zero temperatures [11, 12, 13]. Furthermore, the characterization of quantum critical points is often hindered by unknown order parameters, particularly in systems undergoing multiple non-conventional phase transitions [1, 3, 14]. Traditional approaches, such as Landau’s theory, which links order parameters with symmetry changes, become less effective in cases like topological phase transitions or phases without classical symmetry breaking [15, 16, 17].
@@
In general, approximating the ground state of a system is known to be hard, even for a quantum computer [10]. Nonetheless, variational quantum approaches offer a practical alternative for creating states that capture correctly the physics of the ground state in some cases [18, 19]. These methods have become increasingly important in the study of quantum phase transitions [20, 21, 22, 23, 24, 25, 26, 27]. Particularly relevant and noteworthy are the works of Dreyer et al. [20] and Bosse et al. [24]. In Ref. [20] the authors utilized variational quantum optimization on the transverse-field Ising chain, revealing an intriguing behavior of the order parameter, i.e. the transverse magnetization: it exhibits scaling collapse with respect to the circuit depth exclusively on one side of the phase transition, illustrating distinct behaviors across the critical point. Bosse et al. [24] explored the utility of the variational quantum eigensolver for learning phase diagrams of quantum systems, focusing on the use of both the change of the energy of the output states during optimization and a variety of traditional order parameters. This study underscores the potential of (low-fidelity) variationally optimized quantum states to accurately predict phase transitions across a variety of quantum models, including both one-dimensional and two-dimensional spin and fermion systems. However, it is possible that even with access to accurate polynomially sized quantum circuits, the classical optimization landscape may become filled with locally trapped states [28]. Although these states may effectively reflect the physical properties of various phases, accurately identifying the specific order parameter that defines the phase they represent remains a challenging task.
@@
At the same time, advancements in machine learning (ML) have opened new ways for analysing quantum many-body problems [29, 30, 31]. Notably, machine learning techniques—especially those under the framework of unsupervised learning—have been extensively employed to detect quantum phase transitions [32, 33, 34, 35, 36, 37, 38, 39]. Studies such as those by Huang et al. [39] and Lewis et al. [40] have further theoretically highlighted the efficacy of these algorithms in learning unknown properties of quantum systems’ ground states, significantly within identical phases. This capability to distinguish ground-state data from different phases makes machine learning a potent tool for detecting phase transitions by incrementally adjusting the Hamiltonian parameter and modifying the dataset. This philosophy leverages machine learning’s differential performance across datasets to accurately identify critical phase changes. Despite these innovations, most studies rely on data from exact ground states calculated via methods like exact diagonalization or quantum Monte Carlo simulations, which may limit the analysis of larger or classically intractable systems. Moreover, the determination of the order parameter following training remains ambiguous.
@@
Inspired by these advancements, this work aims to study the use of locally trapped states to characterize phase transitions, without prior knowledge of the associated order parameters. We employ classical machine learning to distill meaningful patterns from these states, thereby learning effective order parameters. Local traps, including local minima, commonly found in the complex energy landscapes of quantum systems, often hinder the efficiency of optimization algorithms by preventing convergence to the global minimum [28]. Nevertheless, they can contain important information about the phase of a quantum system. Fig. 1 provides a schematic overview of our approach. When the optimization deficit is relatively small—that is, the locally trapped states have a non-negligible overlap with the ground state—we show that classical machine learning can successfully detect ground state phase transitions. Indeed, in certain cases, a ground state phase transition affects the spectrum and properties of the low-lying excited states. Under these circumstances, we expect that the machine learning protocol will remain effective, even when the variationally optimized quantum state possesses minimal overlap with the ground state. As quantum phase transitions are well defined only in the thermodynamic limit, applying our approach to finite-size systems can only provide approximate estimates of the critical points. The finite-scaling technique is well-establish method to overcome this issue and to obtain accurate values of critical properties in the thermodynamic limit by extrapolating results obtained within finite-size systems [41, 42]. In this work we show that, under certain assumptions, our algorithm can be employed to implement a variant of this technique in terms of quantum resources, quantified by the depth of the variational quantum circuits, that we term finite-depth extrapolation technique.
@@
In this study, we deploy a quantum optimization-machine learning hybrid algorithm to detect both traditional and topological phase transitions across various quantum systems, including the 1D and 2D transverse-field Ising models (TFIMs) and the extended Su-Schrieffer-Heeger (eSSH) model. Our methodology incorporates linear and deep learning machine learning techniques, namely LASSO [43] and Transformer [44]. We show that our protocol not only precisely identifies critical points but also discerns potential order parameters associated with the phase transitions. Notably, for the 1D TFIM, our approach outperforms traditional magnetization measurements and, through the application of the finite-depth extrapolation technique, our approach determines the critical point of the model’s quantum phase transition with remarkable accuracy despite the maximum fidelity achieved by our shallow variational circuits near this point being only 0.87. In the 2D TFIM, we show that our algorithm accurately predicts phase transitions despite considerable gate noise, with validation provided through empirical experiments on Rigetti’s Ankaa 9Q-1 quantum device. Finally, for the eSSH model, we demonstrate that our method effectively learns an order parameter with minimal finite-size effects, enabling more precise location of phase transitions compared to the traditional finite-size partial-reflection many-body topological invariant [22]. We demonstrate that these learned parameters are also applicable to sketching the phase diagram of related models under varying conditions such as external fields. Crucially, unlike previous studies requiring exact ground states, our methodology operates effectively without such stringent prerequisites.
This paper is structured as follows: Section II presents a comprehensive framework of our protocol, starting with an introduction to the LASSO and Set Transformer algorithms, followed by an exposition of our methodology for detecting phase transitions. In Section III we apply our approach to the 1D and 2D TFIMs and to the eSSH model, sketching their phase diagram, and discussing the learned effective order parameters, including many-body topological invariants [45]. The conclusions and future outlook are discussed in Section IV.
II General Framework
II.1 Machine Learning Preliminaries
LASSO for order parameter selection.
$\mathcal{C}(\bm{\kappa},\lambda)=\left(\frac{1}{4w}\sum_{i=1}^{2w}\left(l_{i}-\kappa_{0}-\sum_{j=1}^{\ell}\kappa_{j}f_{ij}\right)^{2}+\lambda\sum_{j=1}^{\ell}\left|\kappa_{j}\right|\right),$ (1)
$\mathcal{P}_{f}=\{O_{1},O_{2},\ldots,O_{\ell}\},$ (2)
$\bm{f}_{i}=\left(\operatorname{Tr}(O_{1}\rho_{i}),\operatorname{Tr}(O_{2}\rho_{i}),\ldots,\operatorname{Tr}(O_{\ell}\rho_{i})\right),$ (3)
Transformers for complex order parameter synthesis.
The Transformer neural network, developed by Vaswani et al. in 2017, significantly advances the handling of sequential data with its principal component: the self-attention mechanism, also known as scaled dot-product attention [44]. This mechanism enables the Transformer to dynamically prioritize different segments of the input feature vector $\bm{f}_{i}$ based on the relevance of different features, quantified as
where $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, and $\mathbf{V}_{i}$ represent the queries, keys, and values respectively, each derived from the input feature vectors $\bm{f}_{i}$. Here, $d_{k}$ is the dimension of the keys. Queries highlight the current focus within the input, keys facilitate the alignment of these queries with relevant data points, and values convey the substantial data intended for output. The attention formula enables the Transformer to adjust attention across the features dynamically, which enhances its ability to discern complex dependencies. This capability is particularly valuable in detecting quantum phase transitions, as it not only identifies and emphasizes critical observables but also elucidates the intricate dependencies among their expectation values, thereby enabling a comprehensive understanding and synthesis of observables related to the phase transitions.
The multi-head attention concept [44] further extends this capability, enabling the model to dynamically assign different weights of significance to various segments of the input data. This is achieved through the multi-heat attention,
  1. 1.
    Normalization: The input feature vectors $\{\bm{f}_{i}\}$ undergo layer normalization to reduce internal covariate shift, enhancing learning stability.
  2. 2.
    Multi-head attention: The normalized feature vectors $\{\tilde{\bm{f}}_{i}\}$ processed with multi-head attention, allowing the model to concurrently analyse multiple features of the input set. Output from this stage includes a dropout step to prevent overfitting.
  3. 3.
  4. 4.
We utilize the Adam optimizer [47] with an appropriate learning rate to optimize parameters within this framework, aiming to minimize the mean-square loss between the predicted phase labels and the assigned ones. Through its comprehensive handling of feature interdependencies, the Set Transformer exhibits better performance in detecting phase transitions with complex and non-local order parameters. The architectural details and operational functionalities of our Set Transformer-based regression model are schematically depicted in Fig. 2. A demonstrative code for the Set Transformer architecture utilized in our experiments is provided in Appendix A.
In this work, we employ a LASSO-based algorithm to detect phase transitions between the ferromagnetic and paramagnetic phases within 1D and 2D TFIMs in Sections III.1 and III.2, respectively, and utilize a Transformer-based algorithm for identifying topological phase transitions among trivial, symmetry-broken, and topological phases in the eSSH model in Section III.3.
II.2 Methodology
Quantum local traps
Subsequently, we introduce a variational quantum circuit $U(\bm{\gamma},\bm{\beta};g)$ and apply the Fourier series method for its optimization, a strategy described by Zhou et al. [48]. For a $p$-layer quantum circuit, we express certain rotation angles within the $j$-th layer as
Here, $\bm{\gamma}$ and $\bm{\beta}$ are indirectly optimized via the direct optimization of the vectors $\bm{\zeta}$ and $\bm{\eta}$. For the purposes of this study, we have selected $M=4$ as the degree of these polynomials. Leveraging the Fourier strategy alongside the polynomial representation enables the simultaneous global optimization of $\bm{\gamma}$ and $\bm{\beta}$ across various values of $g$. This approach facilitates the minimization of the energy function sum
$\mathcal{E}_{\text{opt}}=\sum_{g\in\mathcal{G}_{\text{opt}}}\operatorname{Tr}\left(\mathcal{H}(g)\rho(\bm{\zeta},\bm{\eta};g)\right),$ (10)
where $\rho(\bm{\zeta},\bm{\eta};g)=U(\bm{\zeta},\bm{\eta};g)\rho_{\text{in}}U^{\dagger}(\bm{\zeta},\bm{\eta};g)$ are the optimized states. $\mathcal{E}_{\text{opt}}$ is theoretically lower bounded by the sum of the ground state energies on $\mathcal{G}_{\text{opt}}$, $\mathcal{E}_{\text{opt}}\geq\sum_{g\in\mathcal{G}_{\text{opt}}}E_{\text{gs}}(g)$. Starting with a variational quantum circuit with a single layer ($p=1$), we uniformly sample the initial values of $\bm{\zeta},\bm{\eta}$ from a predefined range. The optimization of $\bm{\zeta},\bm{\eta}$ is conducted through the Broyden–Fletcher–Goldfarb–Shanno algorithm [51, 52, 53, 54], and the outcomes are then used to guide subsequent optimizations for circuits with increasing depth $p+1$. With the increase of circuit depth $p$, the optimization deficit decreases, enhancing the precision of our estimates for the locations of critical points.
Adopting a global optimization approach, in contrast to individual optimizations for each $g$ individually, confers significant benefits. Firstly, this strategy is markedly more efficient, as, post-optimization, parameters corresponding to any given $g$ value can be readily generated, enabling the preparation of the associated state through circuit execution. Furthermore, global optimization ensures stability across closely related $g$ values, such as $g=0.400$ and $g=0.401$, allowing for the generation of states with consistent features. If the optimization were conducted separately for each $g$, even adjacent values of $g$ could potentially settle into different local traps, leading to distinct states. This divergence could result in highly unstable data unsuitable for machine learning analysis.
The quantum optimization landscape is notoriously swamped with local traps [28], presenting significant challenges in finding the global minimum. In this work, the quantum states we obtain from the variational quantum optimization, denoted as $\rho(\bm{\zeta},\bm{\eta};g)$, are characterized as locally trapped states that have higher energy than the global minima for several reasons. Firstly, the Fourier strategy approach utilized, as discussed in Ref. [48], tends to be trapped in local minima when new parameters are extended by appending zero-vectors without integrating random perturbations upon the increase of $p$. Secondly, our approach is limited by the sampling of a single set of initial parameters for the optimization process, constraining the comprehensive exploration of the available parameter space. Lastly, we model the variables ${\tilde{\gamma}_{k}}$ and ${\tilde{\beta}_{k}}$ as low-degree polynomial functions of the Hamiltonian parameter $g$. This choice, while facilitating computational efficiency, limits full optimization for each $g$ value, thereby reducing the chances of finding the global minima.
2:Establish a variational quantum circuit with parameters as polynomial functions of $g$ with coefficients ${\bm{\zeta},\bm{\eta}}$. Denote the output stated as $\rho(\bm{\zeta},\bm{\eta};g)$. $\triangleright$ Initialization
3:Optimize $\{\bm{\zeta},\bm{\eta}\}$ to minimize $\mathcal{E}_{\text{opt}}=\sum_{g\in\mathcal{G}_{\text{opt}}}\operatorname{Tr}\left(\mathcal{H}(g)\rho(\bm{\zeta},\bm{\eta};g)\right)$. $\triangleright$ Finding locally-trapped states
4:for each $g$ in $\mathcal{G}_{\text{det}}$ do
5:     Prepare optimized quantum state $\rho(\bm{\zeta},\bm{\eta};g)$.
6:     Measure Pauli terms in $\mathcal{P}_{f}$ and record the vector of expectation values $\bm{f}(g)$. $\triangleright$ Record feature vectors for ML analysis
7:end for
12:         Assign label $-1$ to $\bm{f}(\tilde{g})$, add $(\bm{f}(\tilde{g}),-1)$ to $\mathcal{D}$. $\triangleright$ Label pre-transition states
13:     end for
15:         Assign label 1 to $\bm{f}(\tilde{g})$, add $(\bm{f}(\tilde{g}),1)$ to $\mathcal{D}$. $\triangleright$ Label post-transition states
16:     end for
19:end for
Classical loss landscape
The second stage of our methodology involves constructing a depiction of the classical machine learning loss function landscape across various magnitudes of the parameter $g$ by employing classical regressors. Our objective is to pinpoint critical points of phase transitions, which can be identified by analysing the valleys in this landscape. The number of these valleys is utilized as an approximation for the total count of phase transitions. To facilitate this exploration, we set up a detection grid, $\mathcal{G}_{\text{det}}$, consisting of a sequence of points $\{\hat{g}_{\text{min}},\hat{g}_{\text{min}}+\delta\hat{g},\hat{g}_{\text{min}}+2\delta\hat{g},\ldots,\hat{g}_{\text{max}}\}$ with $\delta\hat{g}$ being the detection sampling resolution. For each $g$ in $\mathcal{G}_{\text{det}}$, we prepare the trapped quantum state $\rho(\bm{\zeta},\bm{\eta};g)$, then measure and record the feature vector, $\bm{f}(g)$, which comprises the expectation values of selected Pauli operators, represented by the set $\mathcal{P}_{f}$, from this state.
Our technique, termed presupposed phase label regression, effectively delineates phase boundaries using learnable order parameters. The analysis employs a sliding window technique with a predetermined window size, $w\in\mathbb{N}^{+}$, to systematically explore the parameter space. Initially, we create an empty set to record training losses, denoted as $\mathcal{L}=\varnothing$. Subsequently, for each value of $g$ within the recalibrated range $[\hat{g}_{\text{min}}+w\cdot\delta\hat{g},\hat{g}_{\text{max}}-w\cdot\delta\hat{g}]$, the process is as follows: we start with an empty training set $\mathcal{D}=\varnothing$; for each $\tilde{g}$ in the range $\{g-w\cdot\delta\hat{g},\ldots,g-\delta\hat{g}\}$, we assign a label of $-1$ and add the pair $(\bm{f}(\tilde{g}),-1)$ to $\mathcal{D}$; similarly, for each $\tilde{g}$ in $\{g+\delta\hat{g},\ldots,g+w\cdot\delta\hat{g}\}$, we assign a label of 1 and incorporate the pair $(\bm{f}(\tilde{g}),1)$ into $\mathcal{D}$. Following this, we train a supervised learning regressor (such as LASSO or Transformer) using the dataset $\mathcal{D}$, record the training loss, and append this loss to $\mathcal{L}$. The presence of valleys within $\mathcal{L}$ indicates potential phase transitions, as they represent points where the supervised learning model discerns a significant distinction between pre-transition and post-transition states, effectively leveraging learnable order parameters. The selection of the window size $w$ is critical: if too small, the detection may lack stability; conversely, if too large, the detection may lack sensitivity and precision. The selection of $w$ is influenced by the complexity of our model, allowing for adaptive adjustments to secure an optimal window size. Notably, our methodology exhibits significant robustness to variations in hyperparameters, thereby demonstrating its stability across a broad range of computational settings. For more detailed analysis, refer to Appendix B.
The pseudocode of the algorithm is shown in Fig. 3.
Noise robustness
Theorem 1 (Robustness to Quasi-GD Noise)
Let $\bm{\zeta}$ and $\bm{\eta}$ be optimized parameters, with $\rho_{\text{in}}$ representing the density matrix of the input state. For a Hamiltonian parameter $g$, consider $\rho(\bm{\zeta},\bm{\eta};g)$ as the corresponding ideal output state, and $\mathcal{N}_{\bm{\zeta},\bm{\eta};g}(\cdot)$ as the corresponding noisy quantum channel. If for any $g$ in $\mathcal{G}_{\text{det}}$ and any Pauli operator $O$ in $\mathcal{P}_{f}$, the expectation values satisfy
$\operatorname{Tr}\left(O\Lambda_{\epsilon}\left(\rho(\bm{\zeta},\bm{\eta};g)\right)\right)=\operatorname{Tr}\left(O\mathcal{N}_{\bm{\zeta},\bm{\eta};g}\left(\rho_{\text{in}}\right)\right),$ (11)
where $\Lambda_{\epsilon}(\cdot)$ denotes a global depolarizing channel with a fixed noise rate $\epsilon$, then the channel $\mathcal{N}_{\bm{\zeta},\bm{\eta};g}(\cdot)$ is defined as a quasi-global depolarizing channel with respect to $\{\rho(\bm{\zeta},\bm{\eta};g)\}$ and $\mathcal{P}_{f}$. Given these conditions, the machine learning algorithms—LASSO and Transformer—are capable of predicting critical points from noisy quantum data that are consistent with those predicted from ideal quantum data.
This theorem implies that for noise channels which act like global depolarizing channels for $\mathcal{P}_{f}$, the shape of the classical loss function remains the same. In other words, the machine learning subroutine effectively performs automatic quantum error mitigation in detecting phase transitions. The proof of Theorem 1 and further discussions are provided in Appendix C.
II.3 Finite-Depth Extrapolation
$\tilde{g}_{c}(n)=g_{c}(n\rightarrow\infty)+bn^{-\mu},$ (12)
Furthermore, in our work we introduce a novel extrapolation approach, termed finite-depth extrapolation. This method involves executing our algorithm across a spectrum of circuit depths $p$, within a specifically chosen system size $n$. For each $p$, the value of $n$ is chosen to ensure that quantum information propagation is localized without experiencing boundary effects [23]. For instance, in one-dimensional systems with nearest-neighbour interactions where each layer promotes single-site quantum information propagation, we set $n\geq 2p+2$ to prevent boundary effects. This approach allows the expectation values of certain geometrically central local observables to align with those anticipated for an infinitely large system. Therefore, this allows us to focus on the impact of circuit depth rather than system size.
Through the application of machine learning techniques, an effective order parameter is discerned, and the landscape of the loss function provides an estimation of the phase transition critical point across different values of $p$. Interestingly, in specific scenarios it seems that the location of the critical point scales exponentially with $p$ rather than polynomially. This suggests that if a circuit with $n=\text{poly}(p)$ suffices to avoid boundary effects, then the location of the critical point converges to the thermodynamic limit value exponentially with $n$. Observation 1 provides more details, and this assertion is supported by numerical analyses in Sec. III.1.
Observation 1 (Exponential Precision)
Consider an $n$-qubit quantum system with low-dimensional connectivity, structured such that a $p=\mathrm{poly}(n)$-layer circuit effectively effectively avoids boundary effects. If there exists a local observable serving as an order parameter for a phase transition, the utilization of quantum optimization might allow the achievement of an estimate for the critical point with exponential accuracy, denoted as $\exp(n)$. This level of accuracy significantly exceeds the precision, $\mathrm{poly}(n)$, obtainable through direct preparation of the finite system’s ground state.
Application of the finite-depth extrapolation, as described by
$\tilde{g}_{c}(p)=g_{c}(p\rightarrow\infty)+ce^{-\nu p},$ (13)
The rationale for employing exponential fitting in our finite-depth extrapolation is underpinned by two significant observations. Firstly, as identified in Ref. [20], the energy density of the output states from variational quantum optimization exhibits an exponential dependence with respect to the circuit depth $p$. This characteristic facilitates the determination of critical points, which can be discerned as the derivative of the energy with respect to the Hamiltonian parameter. Secondly, as detailed in Sec. III.1, our numerical tests reveal that the infidelity between the states produced by quantum optimization and the true ground states near the infinite-size critical point diminishes exponentially with increasing $p$. Consequently, the discrepancies in the local order parameters are also expected to reduce exponentially with increasing $p$. Therefore, the finite-depth extrapolation technique we utilize is based on this exponential decay in discrepancies:
$|\tilde{g}_{c}(p)-g_{c}|\sim 1/\text{poly}(2^{p}),$ (14)
thus exponentially converging to the infinite-size critical point as the depth increases.
III Applications
In this section, we apply our algorithm to study quantum phase transitions in various quantum models. We begin with the 1D transverse-field Ising model (TFIM) with periodic boundary conditions. This model, ideal for employing our finite-depth extrapolation method, allows us to verify the exponential scaling of Eq. (13) and to test the algorithm’s resilience to shot noise. Subsequently, our focus shifts to the 2D TFIM with open boundaries, explored through both numerical analysis and experiments on Rigetti’s Ankaa 9Q-1 quantum computer, where we assess the algorithm’s high robustness to gate noise. Finally, we examine the extended Su-Schrieffer-Heeger (eSSH) model [59, 45], characterized by its topological phase transitions. We apply the LASSO-based algorithm to the first two models, where phase transitions are identifiable through local order parameters. In contrast, the eSSH model, which lacks straightforward local order parameters, is analyzed using a Set Transformer-based approach. This method is particularly effective at identifying complex, non-local interdependencies and directly learning non-linear order parameters from the data. For each configuration considered, we initiate the optimization process with a single set of starting values for $\bm{\zeta}$ and $\bm{\eta}$, subsequently employing the Broyden–Fletcher–Goldfarb–Shanno algorithm [51, 52, 53, 54] to obtain local minima states. The quantum circuits are simulated using the Pennylane software framework [60], with all loss functions being linearly normalized within the range $[0,1]$.
III.1 1D Transverse-field Ising Model

### v2

@@ Abstract
Understanding quantum phase transitions in physical systems is fundamental to characterize their behavior at low temperatures. Achieving this requires both accessing good approximations to the ground state and identifying order parameters to distinguish different phases. Addressing these challenges, our work introduces a hybrid algorithm that combines quantum optimization with classical machine learning. This approach leverages the capability of near-term quantum computers to prepare locally trapped states through finite optimization. Specifically, we apply LASSO for identifying conventional phase transitions and the Transformer model for topological transitions, utilizing these with a sliding window scan of Hamiltonian parameters to learn appropriate order parameters and locate critical points. We validated the method with numerical simulations and real-hardware experiments on Rigetti’s Ankaa 9Q-1 quantum computer. This protocol provides a framework for investigating quantum phase transitions with shallow circuits, offering enhanced efficiency and, in some settings, higher precision-thus contributing to the broader effort to integrate near-term quantum computing and machine learning.
@@ I Introduction
Quantum phase diagrams are a fundamental tool to characterize the behavior of quantum systems under varying external conditions, such as temperature and magnetic fields [1, 2, 3, 4, 5, 6]. A better understanding of quantum phase transitions finds application in the field of materials science, where it may inform the development of novel materials with unique properties. For instance, understanding superconductor–insulator phase transitions is instrumental in the research and development of high-temperature superconductors [7, 8]. Due to the intrinsic complexity of quantum systems, the accurate determination of the ground-state phase diagram is a formidable challenge [9, 10]. Experimentally achieving and maintaining the ground state is also difficult, requiring precise control of conditions and often near-zero temperatures [11, 12, 13]. Furthermore, the characterization of quantum critical points is often hindered by unknown order parameters, particularly in systems undergoing multiple unconventional phase transitions [1, 3, 14]. Traditional approaches, such as Landau’s theory, which link order parameters with symmetry changes, often fail for non-trivial topological phase transitions or phases without conventional symmetry breaking [15, 16, 17].
@@
In general, approximating the ground state of a system is known to be hard, even for a quantum computer [10]. Nonetheless, variational quantum approaches offer a practical alternative for creating states that capture correctly the physics of the ground state in some cases [18, 19, 20]. These methods have become increasingly important in the study of quantum phase transitions [21, 22, 23, 24, 25, 26, 27, 28]. Particularly relevant and noteworthy are the works of Dreyer et al. [21] and Bosse et al. [25]. In Ref. [21] the authors utilized variational quantum optimization on the transverse-field Ising chain, revealing an intriguing behavior of the order parameter, i.e. the transverse magnetization: it exhibits scaling collapse with respect to the circuit depth exclusively on one side of the phase transition, illustrating distinct behaviors across the critical point. Bosse et al. [25] explored the utility of the variational quantum eigensolver for learning phase diagrams of quantum systems, focusing on the use of both the change of the energy of the output states during optimization and a variety of traditional order parameters. This study underscores the potential of (low-fidelity) variationally optimized quantum states to provide reliable indicators of phase transitions across a variety of quantum models, including both one-dimensional and two-dimensional spin and fermion systems. However, even with polynomially-sized quantum circuits, classical optimization strategies may be hindered by locally trapped states within the optimization landscape [29, 30]. Although these states may effectively reflect the physical properties of various phases, accurately identifying the specific order parameter that defines the phase they represent remains a challenging task.
@@
At the same time, advancements in machine learning (ML) have opened new ways for analysing quantum many-body problems [31, 32, 33]. Notably, machine learning techniques—especially unsupervised learning—have been extensively employed to detect quantum phase transitions [34, 35, 36, 37, 38, 39, 40, 41, 42]. Studies such as those by Huang et al. [41], Lewis et al. [43], and Onorati et al. [44, 45] have further theoretically highlighted the efficacy of these algorithms in learning unknown properties of quantum systems’ ground states, significantly within identical phases, even in the absence of exponentially decaying correlations. This capability to distinguish ground-state data from different phases makes machine learning a potent tool for detecting phase transitions by incrementally adjusting the Hamiltonian parameter and modifying the dataset. This philosophy leverages machine learning’s differential performance across datasets to accurately identify critical phase changes. Despite these innovations, most studies (e.g., refs. [35, 38, 39, 41, 42]) were focusing on data from exact ground states calculated via methods like exact diagonalization or quantum Monte Carlo simulations, which may limit the analysis of larger or classically intractable systems. Moreover, the determination of the order parameter following training remains ambiguous.
@@
Inspired by these advancements, this work aims to study the use of locally trapped states to characterize phase transitions, without prior knowledge of the associated order parameters. We employ classical machine learning to distill meaningful patterns from these states, thereby learning effective order parameters. Local traps, including local minima, commonly found in the complex energy landscapes of quantum systems, often hinder the efficiency of optimization algorithms by preventing convergence to the global minimum [29, 46]. Nevertheless, they can contain important information about the phase of a quantum system. Fig. 1 provides a schematic overview of our approach. When the optimization deficit is relatively small—that is, the locally trapped states have a non-negligible overlap with the ground state—we show that classical machine learning can successfully detect ground-state phase transitions. Indeed, in certain cases, a ground-state phase transition affects the spectrum and properties of the low-lying excited states. Under these circumstances, we expect that the machine learning protocol will remain effective, even when the variationally optimized quantum state possesses minimal overlap with the ground state. As quantum phase transitions are well defined only in the thermodynamic limit, applying our approach to finite-size systems can only provide approximate estimates of the critical points. The finite-size-scaling technique is a well-established method to overcome this issue and to obtain accurate values of critical properties in the thermodynamic limit by extrapolating results obtained within finite-size systems [47, 48]. In this work we show that, under certain assumptions, our algorithm can be employed to implement a variant of this technique in terms of quantum resources, quantified by the depth of the variational quantum circuits, that we term finite-depth extrapolation technique.
@@
In this study, we deploy a quantum optimization-machine learning hybrid algorithm to detect both traditional and topological phase transitions across various quantum systems, including the 1D and 2D transverse-field Ising models (TFIMs) and the extended Su-Schrieffer-Heeger (eSSH) model. Our methodology incorporates linear and deep learning machine learning techniques, namely LASSO [49] and Transformer [50]. In quantum information, LASSO has been employed for tasks like predicting ground state properties [43] and for learning unknown observables [51]. Here, we apply it to extract effective order parameters and accurately determine critical boundaries. Transformer models, while previously utilized, for example, to perform time-domain analysis from quantum Monte Carlo data [52] or serving as ansätze within Monte Carlo frameworks [53], are implemented in our work in a different way and with a distinct purpose: to capture non-trivial phase-transition-related non-local interdependencies within the data arising from locally trapped states from variational quantum circuits. We show that our protocol can identify critical points and discern candidate order parameters associated with the phase transitions. Notably, for the 1D TFIM, our approach outperforms traditional magnetization measurements and, through the application of the finite-depth extrapolation technique, our approach determines the critical point of the model’s quantum phase transition with high accuracy despite the maximum fidelity achieved by our shallow variational circuits near this point being only 0.75. In the 2D TFIM, we show that our algorithm accurately predicts phase transitions despite considerable gate noise, with validation provided through noisy experiments on Rigetti’s Ankaa 9Q-1 quantum device. Finally, for the eSSH model, we demonstrate that our method effectively learns an order parameter with minimal finite-size effects, yielding more precise critical‑point estimates than those from the finite-size partial-reflection many-body topological invariant. [23].
While our demonstrations in some scenarios employ relatively shallow variational circuits and local observables—raising potential concerns about classical simulability in certain regimes [54, 55]—our approach is not restricted to this setting. In practice, polynomial- or even logarithmic-depth circuits may remain classically intractable for large systems, particularly in the presence of long-range interactions or complex connectivity. Rather than making a formal claim of quantum advantage, our primary goal is to show that reliable phase transition detection remains possible even if the variational circuit does not closely approximate the ground state. We discuss the interplay between our method and potential classical simulability, along with other limitations, in the Discussion. Crucially, although many previous studies have focused on data derived from exact ground states for theoretical and numerical analyses, our methodology demonstrates that reliable phase transition detection can be achieved using variationally optimized local minima states. These states capture the essential physics even at shallow circuit depths, despite their deviation from the exact ground state.
II Results
In this section, we apply our algorithm, described in Section IV, to study quantum phase transitions in various quantum models. We begin with the 1D transverse-field Ising model (TFIM) with periodic boundary conditions. This model, particularly well-suited for applying our finite-depth extrapolation method, allows us to verify the exponential scaling of Eq. (38) and to test the algorithm’s resilience to shot noise. Subsequently, our focus shifts to the 2D TFIM with open boundaries, explored through both numerical analysis and experiments on Rigetti’s Ankaa 9Q-1 quantum computer, where we assess the algorithm’s robustness to moderate levels of gate noise in our tested settings. Finally, we examine the non-integrable extended Su-Schrieffer-Heeger (eSSH) model [56, 57], characterized by its topological phase transitions. We apply the LASSO-based algorithm to the first two models, where phase transitions are identifiable through local order parameters. In contrast, the eSSH model, which lacks straightforward local order parameters, is analysed using a Set Transformer-based approach. This method demonstrates effectiveness in at identifying complex, non-local interdependencies and directly learning non-linear order parameters from the data. For each configuration considered, we initiate the optimization process with a single set of starting values for $\boldsymbol{\zeta}$ and $\boldsymbol{\eta}$, subsequently employing the Broyden–Fletcher–Goldfarb–Shanno algorithm [58, 59, 60, 61] to obtain local minima states. The quantum circuits are simulated using the Pennylane software framework [62], with all loss functions being linearly normalized within the range $[0,1]$.
II.1 1D Transverse-field Ising Model